# AlignDistil: Token-Level Language Model Alignment as Adaptive Policy Distillation

### v1

@@ Abstract
Songming Zhang1,2thanks:  Work was done when Songming was interning at Tencent., Xue Zhang1, Tong Zhang2, Bojie Hu2, Yufeng Chen1thanks:  Yufeng Chen is the corresponding author., and Jinan Xu1 1 Beijing Key Lab of Traffic Data Analysis and Mining, Beijing Jiaotong University, Beijing, China, 2 Tencent Inc, China {smzhang22,zhang_xue,chenyf,jaxu}@bjtu.edu.cn
@@ 6 Analysis
Lin et al. (2024) reveal that DPO reward often shows inferior generalization performance than a pure reward model. We also verify this phenomenon in Table 2. Specifically, we calculate the response-level accuracy of different types of rewards on the training set and the test set of UltraFeedback. Table 2 shows a performance gap between the DPO reward and the reward model. By contrast, the contrastive DPO reward not only shows better accuracy than the vanilla DPO reward on training data, but also generalizes better on test data, even surpassing the reward model. Correspondingly, our on-policy AlignDistil with the contrastive DPO reward outperforms the one with the vanilla DPO reward on AlpacaEval 2.0. This performance gain can be attributed to the reverse DPO model, which captures subtle features in low-quality responses. Therefore, the contrastive DPO reward plays a key role in our AlignDistil.
@@ 7 Related Work
Knowledge distillation (KD, Hinton, 2015) is proposed as an essential technique for compressing neural networks. With the emergence and development of LLMs, KD has attracted more attention to reduce the numerous parameters in LLMs. In this context, KD methods are divided into white-box KD (Hinton, 2015) and black-box KD (Kim and Rush, 2016), depending on whether the weight of the teacher model can be obtained. For white-box KD, approaches typically bridge probability distributions Agarwal et al. (2024); Gu et al. (2024); Ko et al. (2024); Zhang et al. (2024) or intermediate features Wang et al. (2020) between the teacher model and the student model. Intuitively, this process transfers sufficient information from the teacher model, thus often used for pre-training small yet powerful LLMs Team et al. (2024); Meta (2024). In contrast, black-box KD is actually more widely used for LLMs as it only requires collecting outputs from the teacher model and supervised fine-tuning the student model Taori et al. (2023); Chiang et al. (2023); Tunstall et al. (2023). Different from these methods, our AlignDistil is derived from RLHF and aims for token-level reward optimization.
@@ 8 Conclusion
Due to the resource limitation, the evaluation of our AlignDistil is limited within small language models (~1.5B). The evaluation on larger models is still under-explored and we leave this in future work.
@@ Appendix C Implementation Details
In this section, we provide the implementation details of baseline methods and our AlignDistil. All our implementation is based on the open-source toolkit OpenRLHF [^5] https://github.com/OpenRLHF/OpenRLHF. Below we list the individual settings for each method in our experiments:

### v2

@@ Abstract
Songming Zhang1,2thanks:  Work was done when Songming was interning at Tencent., Xue Zhang1,2, Tong Zhang3, Bojie Hu3, Yufeng Chen1,2thanks:  Yufeng Chen is the corresponding author., and Jinan Xu1,2 1Key Laboratory of Big Data & Artificial Intelligence in Transportation, (Beijing Jiaotong University), Ministry of Education 2School of Computer Science and Technology, Beijing Jiaotong University, Beijing, China 3 Tencent Inc, China {smzhang22,zhang_xue,chenyf,jaxu}@bjtu.edu.cn
@@ 6 Analysis
Lin et al. (2024) reveal that DPO reward often shows inferior generalization performance than a pure reward model. We also verify this phenomenon in Table 2. Specifically, we calculate the response-level accuracy of different types of rewards on the training set and the test set of UltraFeedback. Table 2 shows a performance gap between the DPO reward and the reward model. By contrast, the contrastive DPO reward not only shows better accuracy than the vanilla DPO reward on training data, but also generalizes better on test data, even surpassing the reward model. Correspondingly, our on-policy AlignDistil with the contrastive DPO reward outperforms the one with the vanilla DPO reward on AlpacaEval 2.0. This performance gain can be attributed to the reverse DPO model, which captures subtle features in low-quality responses and implicitly doubles the trainable parameters in the DPO reward. Therefore, the contrastive DPO reward plays a key role in our AlignDistil.
@@ 6.2 Effect of Adaptive Logit Extrapolation
Methods Win Rates (%) Avg. Reward
SFT 50% 0.55
DPO 85.7% 1.67
PPO 87.7% 2.22
RTO 90.8% 1.98
Off-Policy AlignDistil 92.8% 2.28
On-Policy AlignDistil 92.5% 2.46
Table 4: Win Rates (%) compared to SFT baseline and the average reward scores of different methods on TL;DR.
6.3 Performance on TL;DR
To further evaluate the performance of our AlignDistil, we also conduct experiments using Qwen2.5-1.5B on TL;DR Stiennon et al. (2022). For evaluation, we calculate the win rate of each method compared to SFT with Qwen2.5-72B-Instruct as the judge model. Besides, we calculate the average reward of each method by averaging the reward scores from the reward model in PPO training. More details on training and evaluation are listed in Appendix C.2. As shown in Table 4, all alignment methods significantly outperform vanilla SFT. On this basis, PPO further outperforms DPO on both win rates and reward, showcasing the benefits of RL training. RTO shows a better win rate than PPO, while the reward is lower than that of PPO. The reason we conjecture is that PPO directly optimizes this reward, while RTO optimizes the DPO reward. Despite this, both our off- and on-policy AlignDistil surpass PPO and RTO on both metrics, which demonstrates the consistently exceptional performance of our method on different benchmarks.
@@ 7 Related Work
Knowledge distillation (KD, Hinton, 2015) is proposed as an essential technique for compressing neural networks. With the emergence and development of LLMs, KD has attracted more attention to reduce the numerous parameters in LLMs. In this context, KD methods are divided into white-box KD (Hinton, 2015) and black-box KD (Kim and Rush, 2016), depending on whether the weight of the teacher model can be obtained. For white-box KD, approaches typically bridge probability distributions Agarwal et al. (2024); Gu et al. (2024); Ko et al. (2024); Zhang et al. (2024, 2025) or intermediate features Wang et al. (2020) between the teacher model and the student model. Intuitively, this process transfers sufficient information from the teacher model, thus often used for pre-training small yet powerful LLMs Team et al. (2024); Meta (2024). In contrast, black-box KD is actually more widely used for LLMs as it only requires collecting outputs from the teacher model and supervised fine-tuning the student model Taori et al. (2023); Chiang et al. (2023); Tunstall et al. (2023). Different from these methods, our AlignDistil is derived from RLHF and aims for token-level reward optimization.
@@ 8 Conclusion
Due to the resource limitation, the evaluation of our AlignDistil is limited to small language models (~1.5B). The evaluation of AlignDistil on larger models is still under-explored, and we leave this for future work.
Acknowledgments
The research work described in this paper has been supported by the Fundamental Research Funds for the Central Universities (2024JBZY019) and the National Nature Science Foundation of China (No. 62476023, 61976016, 62376019), and the authors would like to thank the anonymous reviewers for their valuable comments and suggestions to improve this paper.
@@ Appendix C Implementation Details
In this section, we provide details on the implementation of baseline methods and our AlignDistil. All our implementation is based on the open-source toolkit OpenRLHF [^5] https://github.com/OpenRLHF/OpenRLHF.
C.1 UltraFeedback
Below, we list the individual settings for each method on UltraFeedback:
@@ C.1 UltraFeedback
C.2 TL;DR
Here, we provide the implementation details of the experiments on TL;DR.
Training.
We start from the base model Qwen2.5-1.5B-Base and first conduct SFT on the SFT version of TL;DR [^6] https://huggingface.co/datasets/trl-lib/tldr. Based on the checkpoint after SFT, we further apply LLM alignment algorithms, including DPO, PPO, RTO and our AlignDistil using the preference dataset of TL;DR [^7] https://huggingface.co/datasets/trl-lib/tldr-preference. The configurations of all training procedures are listed below:
  • SFT: We train the Qwen2.5-1.5B model on the SFT dataset of TL;DR dataset for 1 epoch, with a batch size of 128 and a learning rate of 2e-5.
  • Reward Modeling: Based on the checkpoint after SFT, we train a reward model on the preference data of TL;DR dataset for 1 epoch, with a batch size of 128 and a learning rate of 1e-6. This model is used for both PPO training and evaluation.
  • DPO: In DPO training, $\beta$ is set to 0.1. The batch size and the learning rate are set to 128 and 1e-6 for DPO and all the following methods.
  • PPO: The learning rate of the critic model is set to 9e-6 and the KL coefficient is set to 0.07. Other settings follow the default settings in OpenRLHF.
  • RTO: $\beta$ is set to 0.1 and $\beta_{2}$ is set to 0.07. Other settings are kept the same with PPO.
  • Off-Policy AlignDistil: $r$ is set to 1.
  • Off-Policy AlignDistil: $r$ is set to 2.
Evaluation.
During evaluation, we randomly sample 1000 items of data from the validation set of TL;DR-preference. We use Qwen2.5-72B-Instruct to judge the win rates of different methods against SFT. The prompt for judgement follows Ahmadian et al. (2024). Moreover, we leverage the reward model trained for PPO to measure the average reward of the models after different methods.